AB bisects CD at E. If CE = 2 1/4 in., find CD


1
Expert's answer
2021-09-13T00:26:13-0400

"\text{Bisection is the division of a line into 2 equal parts. From our question, we have that }\\
\text{AB bisects CD at E. Hence if CE = $2\frac{1}{4} in.$, we have that CD = $2.2\frac{1}{4}$ = $4\frac{1}{2}in.$}"